OpenStudy (anonymous):

Can you please help me with these problems? Find the product. 4/5 x 8 15 x 1/3 2/3 x 12 5/10 x 100 7/8 x 35 1/8 x 40 1/5 x 200 2/3 x 300

OpenStudy (anonymous):

1. 32/5

OpenStudy (anonymous):

2. 5

OpenStudy (anonymous):

3. 8

OpenStudy (anonymous):

4.50

OpenStudy (anonymous):

Can you please explain how you are doing this?

OpenStudy (anonymous):

1. 4/5*8/1=32/5

OpenStudy (anonymous):

you got it

jhonyy9 (jhonyy9):

32/5 15/3 =5 24/3 =8 500/10 =50 245/8 40/8 =5 200/5 = 40 600/3 = 200